Assume that the multiplier has a value of 3. Now assume
that the government decides to increase aggregate
demand in an attempt to reduce unemployment. It raises
government expenditure by £100 million with no increase
in taxes. Firms, anticipating a rise in their sales, increase
investment by £200 million, of which £50 million consists
of purchases of foreign machinery. By how much will GDP
rise? (Assume that nothing else changes.)
1
Expert's answer
2019-04-25T09:56:32-0400

The GDP multiplier (government expenditure multiplier) is a coefficient that determines the ratio between the initial increase in government spending and the final impact of this increase on the economy.

How does an increase in government spending affect the economy? The growth of public procurement leads to an increase in production and a corresponding increase in income. Income growth leads to a further increase in demand (consumption), which causes a further increase in production, etc. And the end result of GDP growth is greater than the initial shift caused by the growth of government spending. This is the multiplier effect.

The multiplier depends on the propensity to consume.

And this means that each pound spent by the state will cause an increase in GDP equal to 3 pounds.

Hence, by increasing government spending by 100 million, GDP will increase by 300 million.

Economic Keynesian theory suggests the existence of a multiplier effect on investment. The meaning of this phenomenon is that the investment recovery is capable of providing a significant increase in gross domestic product.

So, 200-50 = 150.

150 * 3 = 450.

With an increase in investment (only domestically; purchase of foreign machinery will increase the GDP of foreign country) by 150 million, GDP will increase by 450 million.

Consequently, the cumulative GDP growth will be 300 + 450 = 750 million pounds

Answer: by 750 million pounds